A small business needs to purchase PAD system pattern-making (PAD) software. PAD charges $6,450 for a single-seat V6.0 lifetime license, plus a $1,500 per year maintenance fee for that license. The small business estimates that it will make an additional $2,790 per year in revenues due to the purchase of this software. How many years will it take for the small business to break-even on the purchase given these cash flows?
Answer #1

Answer

break-even is at total revenue is equal to the total cost

n=number of years

total cost =fixed cost + variable cost

=6450+n* 1500

total revenue =n*2790

breakeven is:

2790n=6450+1500n

1290n=6450

n=5

it will take 5 years to break even.